OpenStudy (anonymous):

Help~! 3/4(m+4) > m+3

OpenStudy (anonymous):

no as 3/4(m+4) =3/4(m)+3/4*(4) =3/4(m)+3<m+3 as 3/4 m < m

OpenStudy (anonymous):

Thanks